# EnKF-C user guide version 2.32.0

### v11

@@ 2.7.3 Adaptive moderation of observations
One of the standard QC procedures in DA is the so called background check, when an observation is compared with the forecast and discarded if the innovation magnitude exceeds some specified threshold. The downside of this approach is that it can not distinguish between situations of an outlier, big model error (e.g. because of an error in forcing), or model divergence. While one probably would like to discard an outlier, it is usually desirable to make use of valid observations, although, perhaps, with a reduced impact, to avoid “over-stressing” the model. In EnKF-C this is achieved by adaptive moderation of the observation impact by restricting the magnitude of the increment from a given observation in observation space by $K$ times magnitude of the spread of the forecast ensemble (Sakov and Sandery, 2017).
@@
where $\sigma_{obs}$ is the observation error standard deviation, $\sigma_{f}$ – forecast ensemble spread, $d$ – innovation, and $K$ – the so called K-factor defined in the main parameter file (sec. 2.4.1). Tests with small models show that setting the $K\geq 2$ makes a marginal impact (if any) on performance of weakly suboptimal systems, while still can be quite beneficial in stressful situations.
@@ 2.7.4 Moderation of spread reduction
The moderating parameter $\alpha\in(0,1]$ specified in the main parameter file via the entry ALPHA allows one to reduce the contraction of ensemble during assimilation, while leaving the increment unchanged (“relaxation to prior spread”, Zhang et al. 2004, eq. 5). It modifies the right multiplied ensemble transform matrix as
@@ 2.7.5 Innovation statistics
In its course CALC calculates some basic innovation statistics: number of observations, mean absolute forecast innovation, mean absolute analysis innovation, mean forecast innovation, mean analysis innovation, mean forecast ensemble spread, and mean analysis ensemble spread. This statistics is provided for each region defined in the main parameter file (sec. 2.4.1), as well as for each time slot defined for asynchronous DA, and for each instrument. By default, EnKF-C defines one statistical region “Global” with extent $[x_{1},x_{2}]=[-999,999],[y_{1},y_{2}]=[-999,999]$.
@@
       printing observation statistics:      region obs.type   # obs.  |for.inn.| |an.inn.|   for.inn.   an.inn.  for.spread  an.spread      ------------------------------------------------------------------------------------------      Tasman             SLA          3003    0.067      0.038      0.033      0.012      0.035      0.025              -4           712    0.058      0.038      0.035      0.013      0.028      0.021              -3           785    0.093      0.040      0.060      0.019      0.052      0.034              -2           700    0.062      0.043      0.030      0.016      0.027      0.021              -1           668    0.049      0.031      0.017      0.004      0.028      0.021               0           138    0.078      0.033     -0.043     -0.016      0.045      0.029               j1         1323    0.070      0.033      0.041      0.016      0.037      0.024               n1          876    0.073      0.042      0.052      0.025      0.036      0.026               g1          785    0.054      0.042     -0.004     -0.009      0.029      0.024               N/A          19    0.101      0.037      0.097      0.031      0.059      0.036             SST          9316    0.346      0.174     -0.215     -0.094      0.358      0.254              -4          2946    0.327      0.166     -0.236     -0.092      0.342      0.245              -3          2733    0.368      0.183     -0.270     -0.133      0.362      0.256              -2          2560    0.352      0.169     -0.167     -0.057      0.370      0.262              -1           580    0.342      0.191     -0.148     -0.093      0.414      0.291               0           497    0.305      0.182     -0.126     -0.075      0.307      0.225               AVHRR      9316    0.346      0.174     -0.215     -0.094      0.358      0.254             TEM           768    0.581      0.365     -0.245     -0.151      0.320      0.251               ARGO        768    0.581      0.365     -0.245     -0.151      0.320      0.251               0-50m       125    0.418      0.230      0.049      0.027      0.365      0.281               50-500m     451    0.678      0.403     -0.266     -0.141      0.360      0.278               >500m       192    0.458      0.365     -0.387     -0.291      0.196      0.170             SAL           768    0.079      0.060      0.014      0.019      0.033      0.028               ARGO        768    0.079      0.060      0.014      0.019      0.033      0.028               0-50m       125    0.079      0.063      0.031      0.035      0.034      0.030               50-500m     451    0.092      0.067      0.026      0.032      0.039      0.032               >500m       192    0.048      0.041     -0.027     -0.021      0.018      0.016     This excerpt shows innovation statistics for the region “Tasman”. It contains sections for SST, SLA and TEM observations. The summary statistics for each observation type is shown at the top of each section; then statistics for days -4, -3, -2, -1 and 0 of a 5-day DAW are shown for the two asynchronous types, SST and SLA. (More generally, the numbering of time intervals corresponds to their positions relative to the analysis time. For more details see sec. 2.6.3.) After that, statistics for particular instruments is shown; “N/A” corresponds to superobservations resulted from merging observations from two or more instruments. (From v1.115.0 there is no superobing across different instruments by default.) For subsurface temperature also statistics for shallow (0–50 m), deep ($>$500 m), and intermediate (50–500 m) observations is given.
@@ 2.7.6 Impact of observations
where $\mathrm{tr}(\cdot)$ is the trace function. The values of these metrics for each local analysis, calculated both for all observations and for observations of each type only, are written to file enkf_diag.nc. Note that the in EnKF-C DFS and SRF are calculated from the above expressions and represent theoretical values for the EnKF analysis; they coincide with the actual DFS and SRF values only for the ETKF, but not for the DEnKF, which is an approximation of the KF (and indeed not for the EnOI, which is not even an approximation). Also note that although the partial DFS and SRF for particular observation types are useful indicators of the impact of the corresponding observations, they are not fully consistent in the sense that these impacts are not independent of each other.
In the EnKF context DFS is a useful indicator of potential rank problems. Normally, it should not exceed a fraction (a half, or better, a quarter) of the ensemble size per the characteristic time of the error growth. SRF shows the “strength” of DA. “Strong” DA implies a close to optimal system, which indeed never happens in practice. Therefore, ideally, SRF should be small (below 1, on average).
@@ 2.7.8 Domains
By default, all local observations for a given grid node contribute to the corresponding ensemble transform. Sometimes it is desirable to disconnect observations of certain type from contributing to transforms on particular grids. For example, it may be desirable in climate systems to disregard observations of the sea surface height in updating the atmospheric variables. The concept of “domains” introduced in v1.89.0 provides a mechanism for handling such situations within a single analysis. It works as follows. Each grid can be associated with a certain domain via the optional entry DOMAIN in the grid parameter file. For example, in a climate model one can have domains “Ocean” and “Atmosphere”. Then entry DOMAINS in the observation types parameter file can list domains observations of this type are visible from. By default, observations of any type are visible from all grids.
@@
$f_{0}(\frac{x}{2})=\left\{\begin{array}[]{ll}1-\frac{5}{3}x^{2}+\frac{5}{8}x^{3}+\frac{1}{2}x^{4}-\frac{1}{4}x^{5},&0\leq x\leq 1,\\ -\frac{2}{3}x^{-1}+4-5x+\frac{5}{3}x^{2}+\frac{5}{8}x^{3}-\frac{1}{2}x^{4}+\frac{1}{12}x^{5},&1<x\leq 2,\\ 0,&2<x.\end{array}\right.$
@@ 2.8 UPDATE
     >./bin/enkf_update    Usage: enkf_update <prm file> [<options>]    Options:    --allow-logspace-with-static-ens        confirm that static ensemble is conditioned for using log space    --calculate-spread        calculate forecast and analysis ensemble spread and write to spread.nc    --calculate-spread-only        calculate forecast and analysis ensemble spread only and write to spread.nc    --calculate-forecast-spread        calculate forecast ensemble spread and write to spread.nc    --calculate-forecast-spread-only        calculate forecast ensemble spread only and write to spread.nc    --describe-prm-format [main|model|grid]        describe format of a parameter file and exit    --direct-write        write fields directly to the output file (default: write to tiles first)    --joint-output        append analyses to forecast files (default: write to separate files)    --no-fields-write        do not write analysis fields (only diagnostic data)    --no-update        exclude tasks that require ensemble update    --output-increment        output analysis increment (default: output analysis)    --write-inflation        write adaptive inflation magnitudes to inflation.nc    --version        print version and exit
The option --joint-output tells UPDATE to append analyses to the corresponding forecast files, using new variable names constructed by concatenating the forecast variable names and suffix _an. By default the analyses are written to separate files, using the same variable names as the forecast files, but with an extra suffix .analysis or .increment added to the file name, depending on whether the analysis or increment is written.
By default, UPDATE first writes each updated horizontal field of the model to a separate file (referred to here as a tile), and then concatenates these fields into analysis files. This approach is somewhat less effective than direct writing to analysis files (without intermediate tiles), but, unfortunately, the direct writing is generally not reliable due to parallel I/O issues with NetCDF. Note that in some cases it proved to be possible to obtain robust performance with direct write using “classic” or “64-bit-offset” NetCDF formats.
@@ 2.9 Hybrid covariance
When the method is set to “hybrid”, the forecast ensemble spread written in the innovation statistics summary at the end of CALC is that of the combined ensemble (1.54), but the analysis innovation spread is that of the dynamic ensemble only. The forecast and analysis ensemble spread fields written by UPDATE when specifying options --calculate-spread, --calculate-forecast-spread, and --calculate-forecast-spread-only are calculated using the dynamic ensemble anomalies only. In contrast, the vertical correlations (UPDATE options --calculate-vertical-correlations and --calculate-vertical-correlations-only) are calculated using the full ensemble.
@@ 2.10 Ensemble diagnostics
In version 2.27.0 a new binary enkf_diag has been added to EnKF-C. It aims at calculating various ensemble diagnostics outside the context of DA cycling. It has the following usage and options:
     >./bin/enkf_diag    Usage: enkf_diag <prm file> [<options>]    Options:    --calculate-spread        calculate ensemble spread and write to spread.nc    --calculate-vertical-correlations        calculate correlation coefficients between surface and other layers of        3D variables and write to vcorr.nc    --calculate-vertical-correlations-with <varname1> <layer1>      [<varname2> <layer2>] [...]        calculate correlation coefficients between specified field (a layer of        a variable) and all other fields on the same horizontal grid and        write to vcorr-<varname>-<layer>.nc    --calculate-vertical-covariances-with <varname1> <layer1>      [<varname2> <layer2>] [...]        calculate covariances between specified field and all other fields        on the same horizontal grid and write to vcov-<varname>-<layer>.nc    --version        print version and exit
@@ 2.11 DA tuning
    R-factors;
    inflation magnitudes;
    localisation radii;
    K-factor.
@@
Localisation radius is defined by the entry LOCRAD in the parameter file. Specifically, this entry defines the support localisation radius (in km). This is different to the “effective” localisation radius, which is defined sometimes as $e^{1/2}\approx 1.65$ folding distance. For the Gaspary and Cohn’s taper function used in EnKF-C the effective radius is approximately 3.5 times smaller than the support radius.
@@ 2.12 Point logs
     netcdf pointlog_156.$0$  dimensions:          m1 = 96 ;          m2 = 96 ;          p-0 = 1473 ;          p-1 = 1473 ;          nk-0 = 2 ;          nk-1 = 2 ;  variables:          int obs_ids(p) ;          float lcoeffs(p) ;          float lon(p) ;          float lat(p) ;          float depth(p) ;          float obs_val(p) ;          float obs_estd(p) ;          float obs_fij0(p) ;          float obs_fij1(p) ;          float obs_fij2(p) ;          float obs_fk(p) ;          int obs_type(p) ;                  obs_type:SLA = 0 ;                  obs_type:RFACTOR_SLA = 2. ;                  obs_type:LOCRAD_SLA = 200. ;                  obs_type:WEIGHT_SLA = 1. ;                  obs_type:GRIDID_SLA = 0 ;                  obs_type:SST = 1 ;                  obs_type:RFACTOR_SST = 4. ;                  obs_type:LOCRAD_SST = 200. ;                  obs_type:WEIGHT_SST = 1. ;                  obs_type:GRIDID_SST = 0 ;                  obs_type:TEM = 2 ;                  obs_type:RFACTOR_TEM = 8. ;                  obs_type:LOCRAD_TEM = 800. ;                  obs_type:WEIGHT_TEM = 1. ;                  obs_type:GRIDID_TEM = 0 ;                  obs_type:SAL = 3 ;                  obs_type:RFACTOR_SAL = 8. ;                  obs_type:LOCRAD_SAL = 800. ;                  obs_type:WEIGHT_SAL = 1. ;                  obs_type:GRIDID_SAL = 0 ;          int obs_inst(p) ;                  obs_inst:j1 = 0 ;                  obs_inst:n1 = 1 ;                  obs_inst:ESACCI = 2 ;                  obs_inst:WindSat = 3 ;                  obs_inst:WMO851 = 4 ;                  obs_inst:WMO846 = 5 ;                  obs_inst:WMO999 = 6 ;          float obs_time(p) ;                  obs_time:units = "days from days from 6565.5 days since 1990-01-01" ;          int grid-0 ;                  grid-0:id = 0 ;                  grid-0:name = "t-grid" ;                  grid-0:domain = "ALL" ;                  grid-0:fi = 49.5 ;                  grid-0:fj = 49.5 ;                  grid-0:nk = 2 ;                  grid-0:model_depth = 4642.25f ;          int grid-1 ;                  grid-1:id = 1 ;                  grid-1:name = "c-grid" ;                  grid-1:domain = "ALL" ;                  grid-1:fi = 48.9999694824219 ;                  grid-1:fj = 49.0000076293945 ;                  grid-1:nk = 2 ;                  grid-1:model_depth = NaNf ;          float s-0(p-0) ;          float S-0(m2, p-0) ;          double w-0(m2) ;                  w-0:long_name = "ensemble coefficients for location (fi,fj)=(49.500,49.500) on grid 0 (\"t-grid\")" ;          double T-0(m1, m2) ;                  T-0:long_name = "ensemble anomalies transform for location (fi,fj)=(49.500,49.500) on grid 0 (\"t-grid\")" ;          float s-1(p-1) ;          float S-1(m2, p-1) ;          double w-1(m2) ;                  w-1:long_name = "ensemble coefficients for location (fi,fj)=(49.000,49.000) on grid 1 (\"c-grid\")" ;          double T-1(m1, m2) ;                  T-1:long_name = "ensemble anomalies transform for location (fi,fj)=(49.000,49.000) on grid 1 (\"c-grid\")" ;          float eta_t(m2) ;          float eta_t_an(m2) ;                  eta_t_an:gridid = 0 ;                  eta_t_an:INFLATION = 1.1f, 1.f ;          float temp(nk-0, m2) ;                  temp:gridid = 0 ;                  temp:INFLATION = 1.1f, 1.f ;          float temp_an(nk-0, m2) ;          float salt(nk-0, m2) ;                  salt:gridid = 0 ;                  salt:INFLATION = 1.1f, 1.f ;          float salt_an(nk-0, m2) ;          float u(nk-1, m2) ;                  u:gridid = 1 ;                  u:INFLATION = 1.1f, 1.f ;          float u_an(nk-1, m2) ;          float v(nk-1, m2) ;                  v:gridid = 1 ;                  v:INFLATION = 1.1f, 1.f ;          float v_an(nk-1, m2) ;    // global attributes:                  :version = "2.21.3" ;                  :lon = 156. ;                  :lat = -32. ;                  :MODE = "EnKF" ;                  :SCHEME = "DEnKF" ;                  :ALPHA = 1. ;                  :ngrids = 2 ;                  :output = "analysis" ;  }

### v12

@@ 2.7.3 Adaptive moderation of observations
One of the standard QC procedures in DA is the so called background check, when an observation is compared with the forecast and discarded if the innovation magnitude exceeds some threshold. The downside of this approach is that it can not distinguish between situations of an outlier, big model error (e.g. because of an error in forcing), or model divergence. While one probably would like to discard an outlier, it is usually desirable to make use of valid observations, although, perhaps, with a reduced impact, to avoid “over-stressing” the model. In EnKF-C this is achieved by adaptive moderation of the observation impact by restricting the magnitude of the increment from a given observation in observation space by $K$ times magnitude of the forecast ensemble spread (Sakov and Sandery, 2017).
@@
where $\sigma_{obs}$ is the observation error standard deviation, $\sigma_{f}$ – forecast ensemble spread, $d$ – innovation, and $K$ – the so called K-factor defined in the main parameter file (sec. 2.4.1). Tests with small models show that setting the $K\geq 2$ makes a marginal impact (if any) on performance of weakly suboptimal systems, while it still can be quite beneficial in situations with large innovations.
@@ 2.7.4 Moderation of spread reduction
The moderating parameter $\alpha\in(0,1]$ specified in the main parameter file via the entry ALPHA allows one to reduce the contraction of ensemble during assimilation, while leaving the increment unchanged (“relaxation to prior spread”, RPS, Zhang et al. 2004, eq. 5). It modifies the right multiplied ensemble transform matrix as
@@ 2.7.5 Innovation statistics
In its course CALC calculates some basic innovation statistics: number of observations, mean absolute forecast innovation, mean absolute analysis innovation, mean forecast innovation, mean analysis innovation, mean forecast ensemble spread, and mean analysis ensemble spread. This statistics is provided for each region defined in the main parameter file (sec. 2.4.1), as well as for each time slot defined for asynchronous DA, and for each instrument. By default, EnKF-C defines one statistical region “Global” with extent $x_{1},x_{2}]=[-999,999],[y_{1},y_{2}]=[-999,999]$.
@@
       printing observation statistics:     region obs.type # obs. |for.inn.| |an.inn.| for.inn. an.inn. for.spread an.spread     ------------------------------------------------------------------------------------------     Tasman     SLA 3003 0.067 0.038 0.033 0.012 0.035 0.025     -4 712 0.058 0.038 0.035 0.013 0.028 0.021     -3 785 0.093 0.040 0.060 0.019 0.052 0.034     -2 700 0.062 0.043 0.030 0.016 0.027 0.021     -1 668 0.049 0.031 0.017 0.004 0.028 0.021     0 138 0.078 0.033 -0.043 -0.016 0.045 0.029     j1 1323 0.070 0.033 0.041 0.016 0.037 0.024     n1 876 0.073 0.042 0.052 0.025 0.036 0.026     g1 785 0.054 0.042 -0.004 -0.009 0.029 0.024     N/A 19 0.101 0.037 0.097 0.031 0.059 0.036     SST 9316 0.346 0.174 -0.215 -0.094 0.358 0.254     -4 2946 0.327 0.166 -0.236 -0.092 0.342 0.245     -3 2733 0.368 0.183 -0.270 -0.133 0.362 0.256     -2 2560 0.352 0.169 -0.167 -0.057 0.370 0.262     -1 580 0.342 0.191 -0.148 -0.093 0.414 0.291     0 497 0.305 0.182 -0.126 -0.075 0.307 0.225     AVHRR 9316 0.346 0.174 -0.215 -0.094 0.358 0.254     TEM 768 0.581 0.365 -0.245 -0.151 0.320 0.251     ARGO 768 0.581 0.365 -0.245 -0.151 0.320 0.251     0-50m 125 0.418 0.230 0.049 0.027 0.365 0.281     50-500m 451 0.678 0.403 -0.266 -0.141 0.360 0.278     >500m 192 0.458 0.365 -0.387 -0.291 0.196 0.170     SAL 768 0.079 0.060 0.014 0.019 0.033 0.028     ARGO 768 0.079 0.060 0.014 0.019 0.033 0.028     0-50m 125 0.079 0.063 0.031 0.035 0.034 0.030     50-500m 451 0.092 0.067 0.026 0.032 0.039 0.032     >500m 192 0.048 0.041 -0.027 -0.021 0.018 0.016      This excerpt shows innovation statistics for the region “Tasman”. It contains sections for SST, SLA and TEM observations. The summary statistics for each observation type is shown at the top of each section; then statistics for days -4, -3, -2, -1 and 0 of a 5-day DAW are shown for the two asynchronous types, SST and SLA. (More generally, the numbering of time intervals corresponds to their positions relative to the analysis time. For more details see sec. 2.6.3.) After that, statistics for particular instruments is shown; “N/A” corresponds to superobservations resulted from merging observations from two or more instruments. (From v1.115.0 there is no superobing across different instruments by default.) For subsurface temperature also statistics for shallow (0–50 m), deep (>>500 m), and intermediate (50–500 m) observations is given.
@@ 2.7.6 Impact of observations
where $\mathrm{tr}(\cdot)$ is the trace function. The values of these metrics for each local analysis, calculated both for all observations and for observations of each type only, are written to file enkf_diag.nc. Note that in EnKF-C DFS and SRF are calculated from the above expressions and represent theoretical values for the EnKF analysis; they coincide with the actual DFS and SRF values only for the ETKF, but not for the DEnKF, which is an approximation of the KF (and indeed not for the EnOI, which is not even an approximation). Also note that although the partial DFS and SRF for particular observation types are useful indicators of the impact of the corresponding observations, they are not fully consistent in the sense that these impacts are not independent of each other.
In the EnKF context DFS is a useful indicator of potential rank problems. Normally, it should not exceed a fraction (a half, or better, a quarter) of the ensemble size per the characteristic time of the error growth. SRF shows the “strength” of DA. “Strong” DA requires a nearly optimal system (otherwise it produces large artefacts and imbalances), which indeed never happens in practice. Therefore, ideally, SRF should be small (below 1, on average).
@@ 2.7.8 Domains
By default, all local observations for a given grid node contribute to the corresponding ensemble transform. Sometimes it is desirable to disconnect observations of certain type from contributing to transforms on particular grids. For example, it may be desirable in climate systems to disregard observations of the sea surface height (SSH) in updating the atmospheric variables. The concept of “domains” introduced in v1.89.0 provides a mechanism for handling such situations within a single analysis. It works as follows. Each grid can be associated with a certain domain via the optional entry DOMAIN in the grid parameter file. For example, in a climate model one can have domains “Ocean” and “Atmosphere”. Then entry DOMAINS in the observation types parameter file can list domains observations of this type are visible from. By default, observations of any type are visible from all grids.
@@
$f_{0}(\frac{x}{2})=\left\{\begin{array}[]{ll}1-\frac{5}{3}x^{2}+\frac{5}{8}x^{3}+\frac{1}{2}x^{4}-\frac{1}{4}x^{5},\quad&0\leq x\leq 1,\\ -\frac{2}{3}x^{-1}+4-5x+\frac{5}{3}x^{2}+\frac{5}{8}x^{3}-\frac{1}{2}x^{4}+\frac{1}{12}x^{5},\quad&1<x\leq 2,\\ 0,\quad&2<x.\end{array}\right$.
@@ 2.8 UPDATE
      >./bin/enkf_update     Usage: enkf_update <prm file> [<options>]     Options:     --allow-logspace-with-static-ens     confirm that static ensemble is conditioned for using log space     --calculate-spread     calculate forecast ensemble spread and write to spread.nc     --describe-prm-format [main|model|grid]     describe format of a parameter file and exit     --direct-write     write fields directly to the output file (default: write to tiles first)     --no-fields-write     do not write analysis fields (only diagnostic data)     --no-update     exclude tasks that require ensemble update     --output-increment     output analysis increment (default: output analysis)     --write-inflation     write capped inflation magnitudes to inflation.nc     --version     print version and exit
The analysed restarts are written to separate files using the same names as the forecast files but with an extra suffix .analysis or .increment (sec. sec:filenames), depending on whether the analysis or analysis increment is written.
By default, UPDATE first writes each updated horizontal field of the model to a separate file, and then concatenates these fields into analysis files. This approach is somewhat less effective than direct writing to analysis files (without intermediate tiles), but, unfortunately, the direct writing is generally not reliable due to parallel I/O issues with NetCDF. Note that in some cases it proved to be possible to obtain robust performance with direct write using “classic” or “64-bit-offset” NetCDF formats.
Updating a horizontal field requires reading the full ensemble into memory. This may require large available memory, particularly with very large models. From version 2.43.0 it is possible to calculate analyses on fractions of a horizontal field, which may provide better scalability. This option is specified by parameter FIELDSPLIT of the main parameter file. When FIELDPLIT is set to a number greater than one, an additional stage of combining outputs on these fractional tiles into full fields is conducted.
@@ 2.9 Hybrid covariance
When the method is set to “hybrid”, the ensemble spread written in the innovation statistics summary at the end of CALC is that of the combined ensemble (1.54). Similarly, the ensemble spread fields calculated during UPDATE by specifying option --calculate-spread is the ensemble spread of the combined ensemble. To calculate spread of dynamic ensemble only one needs to set MODE = EnKF. To calculate innovation statistics with forecast ensemble spread of the dynamic ensemble only one needs to set MODE = EnKF and rerun CALC with option “–forecast-stats-only”.
@@ 2.10 Ensemble diagnostics
Starting from version 2.27.0 a new binary enkf_diag was added to EnKF-C; from v.2.37.0 it is called ens_diag. It aims at calculating various ensemble diagnostics outside the context of DA cycling. It has the following usage and options:
      >./bin/ens_diag     Usage: enks_diag <prm file> [<options>]     Options:     --calculate-spread     calculate ensemble spread and write to spread.nc     --calculate-vertical-correlations     calculate correlation coefficients between surface and other layers of     3D variables and write to vcorr.nc     --calculate-vertical-correlations-with <varname1> <layer1>     [<varname2> <layer2>] [...]     calculate correlation coefficients between specified field (a layer of     a variable) and all other fields on the same horizontal grid and     write to vcorr-<varname>-<layer>.nc     --calculate-vertical-covariances-with <varname1> <layer1>     [<varname2> <layer2>] [...]     calculate covariances between specified field and all other fields     on the same horizontal grid and write to vcov-<varname>-<layer>.nc     --calculate-vertical-sensitivities-with <varname1> <layer1>     [<varname2> <layer2>] [...]     calculate sensitivities between specified field and all other fields     on the same horizontal grid and write to vcov-<varname>-<layer>.nc     --describe-prm-format [main|model|grid]     describe format of a parameter file and exit     --version     print version and exit      While most of the options above are straightforward, we note that the option “–calculate-vertical-sensitivities-with” calculates the ratio $s_{xy}=\sigma_{x,y}/\sigma_{y,y}$, where $\sigma_{x,y}$ is the covariance between $x$ and $y$, and $\sigma_{y,y}$ is the variance of $y$, so that the increment of $x$ is related to the increment of $y$ as $\Delta x=s_{xy}\Delta y$ (see e.g. Anderson, 2003, eq. 5).
Note that ens_diag can work with a much simpler parameter file than other executables in EnKF-C.
@@ 2.11 DA tuning
    R-factors (sec. 2.4.4);
    localisation radii (sec. 2.4.4);
    inflation magnitudes (sec. 2.4.2,2.8.1);
    K-factor (sec. 2.7.3);
    magnitude of relaxation to prior spread (parameter ALPHA, sec. 2.7.4);
    scaling of static covariance (parameter GAMMA, sec. 2.9.
@@
Localisation radius is defined by the entry LOCRAD in the parameter file. Specifically, this entry defines the support localisation radius (in km). This is different to the “effective” localisation radius, which is defined sometimes as $e^{1/2}\approx 1.65$ - folding distance. For the Gaspary and Cohn’s taper function used in EnKF-C the effective radius is approximately 3.5 times smaller than the support radius.
@@ 2.12 Point logs
      netcdf pointlog_156.000\,-32.000 {    dimensions:   m1 = 96 ;   m2 = 96 ;   p = 1440 ;   p-0 = 1440 ;   p-1 = 1440 ;   nk-0 = 2 ;   nk-1 = 2 ;    variables:   int obs_ids(p) ;   float lcoeffs(p) ;   float lon(p) ;   float lat(p) ;   float depth(p) ;   float obs_val(p) ;   float obs_estd(p) ;   float obs_fij0(p) ;   float obs_fij1(p) ;   float obs_fij2(p) ;   float obs_fk(p) ;   int obs_type(p) ;   obs_type:SLA = 0 ;   obs_type:SST = 1 ;   obs_type:TEM = 2 ;   obs_type:SAL = 3 ;   obs_type:RFACTOR_SLA = 2. ;   obs_type:LOCRAD_SLA = 200. ;   obs_type:WEIGHT_SLA = 1. ;   obs_type:GRIDID_SLA = 0 ;   obs_type:RFACTOR_SST = 32. ;   obs_type:LOCRAD_SST = 200. ;   obs_type:WEIGHT_SST = 1. ;   obs_type:GRIDID_SST = 0 ;   obs_type:RFACTOR_TEM = 8. ;   obs_type:LOCRAD_TEM = 800. ;   obs_type:WEIGHT_TEM = 1. ;   obs_type:GRIDID_TEM = 0 ;   obs_type:RFACTOR_SAL = 8. ;   obs_type:LOCRAD_SAL = 800. ;   obs_type:WEIGHT_SAL = 1. ;   obs_type:GRIDID_SAL = 0 ;   int obs_inst(p) ;   obs_inst:j1 = 0 ;   obs_inst:n1 = 1 ;   obs_inst:ESACCI = 2 ;   obs_inst:WindSat = 3 ;   obs_inst:WMO851 = 4 ;   obs_inst:WMO846 = 5 ;   float obs_time(p) ;   obs_time:units = "days from 6565.5 days since 1990-01-01" ;   int grid-0 ;   grid-0:id = 0 ;   grid-0:name = "t-grid" ;   grid-0:domain = "ALL" ;   grid-0:fi = 49.5 ;   grid-0:fj = 49.5 ;   grid-0:nk = 2 ;   grid-0:model_depth = 4642.25f ;   int grid-1 ;   grid-1:id = 1 ;   grid-1:name = "c-grid" ;   grid-1:domain = "ALL" ;   grid-1:fi = 48.9999691741445 ;   grid-1:fj = 49.0000077064579 ;   grid-1:nk = 2 ;   grid-1:model_depth = NaNf ;   float s-0(p-0) ;   float S-0(m2, p-0) ;   double w-0(m2) ;   w-0:long_name = "ensemble coefficients for location (fi,fj)=(49.500,49.500) on grid 0 (\"t-grid\")" ;   double T-0(m1, m2) ;   T-0:long_name = "ensemble anomalies transform for location (fi,fj)=(49.500,49.500) on grid 0 (\"t-grid\")" ;   float s-1(p-1) ;   float S-1(m2, p-1) ;   double w-1(m2) ;   w-1:long_name = "ensemble coefficients for location (fi,fj)=(49.000,49.000) on grid 1 (\"c-grid\")" ;   double T-1(m1, m2) ;   T-1:long_name = "ensemble anomalies transform for location (fi,fj)=(49.000,49.000) on grid 1 (\"c-grid\")" ;   float eta_t(m2) ;   eta_t:gridid = 0 ;   float eta_t_an(m1) ;   eta_t_an:gridid = 0 ;   eta_t_an:INFLATION = 1.1f, 1.f ;   float temp(nk-0, m2) ;   temp:gridid = 0 ;   float temp_an(nk-0, m1) ;   temp_an:gridid = 0 ;   temp_an:INFLATION = 1.1f, 1.f ;   float salt(nk-0, m2) ;   salt:gridid = 0 ;   float salt_an(nk-0, m1) ;   salt_an:gridid = 0 ;   salt_an:INFLATION = 1.1f, 1.f ;   float u(nk-1, m2) ;   u:gridid = 1 ;   float u_an(nk-1, m1) ;   u_an:gridid = 1 ;   u_an:INFLATION = 1.1f, 1.f ;   float v(nk-1, m2) ;   v:gridid = 1 ;   float v_an(nk-1, m1) ;   v_an:gridid = 1 ;   v_an:INFLATION = 1.1f, 1.f ;        // global attributes:   :lon = 156. ;   :lat = -32. ;   :MODE = "EnKF" ;   :SCHEME = "DEnKF" ;   :ALPHA = 1. ;   :ngrids = 2 ;   :output = "analysis" ;   :EnKF-C\ version = "2.43.3" ;   :command = "./enkf_update --calculate-spread --write-inflation enkf.prm" ;   :wdir = "/home/599/pxs599/src/enkf-c/enkf/examples/4" ;    }